OpenStudy (anonymous):

Find the antiderivative of x/(x+2)^0.5

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=x%2F%28x%2B2%29^0.5dx you can view the detailed answer.

OpenStudy (sriram):

x+2=a^2 dx=2a.da hence x/(x+2)^0.5.dx=a^2 - 2 2(a^2 - 2).da integrate and replace a by x ------- . 2a.da = a

OpenStudy (akshay_budhkar):

strawbocide you get it?

OpenStudy (anonymous):

Awesome, I get it now. Thanks